OpenStudy (anonymous):

Find the x- and y-intercept. 3x-y=9

OpenStudy (anonymous):

Don't tell me the answer, show me how to work it out please.

OpenStudy (danjs):

for the x intercept, the Y coordinate there is Zero, find the point (x , 0) by letting y=0 and solve for x

OpenStudy (danjs):

same for the y intercept, let x=0, and solve for y (0 , y)

OpenStudy (anonymous):

how do i solve for x? sorry kinda new to this

OpenStudy (danjs):

Y intercept, Let x=0 3*0 - y = 9 -y = 9 divide both sides by -1 y = -9 the point x=0, y=-9 (0,-9) is the y-intercept

OpenStudy (danjs):

let Y=0 and solve for X, to get the x intercept. 3x - 0 = 9 3x = 9 divide both sides by 3

OpenStudy (anonymous):

x=3 (3,0) the x intercept